A catapult used to hold a stone of mass 500g is extended by 20cm with an applied force F. If the stone leaves with a velocity of 40ms-1 the value of F is


1
Expert's answer
2021-05-18T11:09:25-0400

According to Newton's second law,


"F=ma,\\\space\\
a=\frac{v^2}{2d},\\\space\\
F=\frac{v^2m}{2d}=2000\text{ N}."